OpenStudy (anonymous):

A little help please? ;; (a) Describe the Relationship between the circumference and area of a circle. (b) Explain the difference between the circumference and area formulas.

OpenStudy (dan815):

okay

OpenStudy (anonymous):

A. TT (often written pi) is a mathematical constant whose value is the ratio between circumference and diameter of any circle in a Euclidean space, is the same value as the ratio of area of a circle and square of its radius B area=pi(r^2) circumference=2*pi*r

OpenStudy (dan815):

circumference = 2*pi*r and the area = pi*r^2

OpenStudy (dan815):

if you knew the circumferece of a circle u can figure out the area

OpenStudy (anonymous):

yep I totally agree with dan815

OpenStudy (anonymous):

Well, yeah. They both use pi and the radius for A. 2piR and piR^2

OpenStudy (dan815):

oops forgot the 2^2 in there

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

B?

OpenStudy (dan815):

as r=circumference/2pi r=c/2pi for convenience so area = pi*(c/2pi)^2 =pi*c^2/(2pi)^2 =c^2/(4pi)

OpenStudy (anonymous):

yeah that is answer for b

OpenStudy (anonymous):

Jk

OpenStudy (anonymous):

B tho.

OpenStudy (anonymous):

I mean I know the difference just a bit hard to put it in words.

OpenStudy (dan815):

okay well to start with you know circumference is a lennght when area is length^2

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

By length you mean radius or diameter?

OpenStudy (anonymous):

????

OpenStudy (dan815):

the units of measurement

OpenStudy (dan815):

it means circumference is measured in one dimention just length and the area is measured in 2 dimensions length and width kind of deal

OpenStudy (dan815):

when you say the area is like 50 cm^2 that means 50 of 1 cm by 1 cm squares fit in your area

OpenStudy (anonymous):

oh, yeah.

OpenStudy (dan815):

when you say 50 cm that means a line that is 50 cm long fit in it or 50 lines of 1 cm long together fit in it

OpenStudy (anonymous):

area is squared and c is a line

OpenStudy (dan815):

ya somethinhg like that

OpenStudy (anonymous):

but i dont think thats what theyre trying to ask.

OpenStudy (anonymous):

our teacher never really touched that, so I'm afraid I won't get it correct bc it's not the answer she wants.

OpenStudy (anonymous):

I think the answer she wants is about the actual formula for the 2.